# Learned Hemodynamic Coupling Inference in Resting-State Functional MRI

### v1

@@ V-B3 Connectivity Analysis
Using the procedure in Section IV-E2, we compared the seed-based effective connectivity for raw BOLD (BOLD), deconvolution with a spatially fixed canonical response (Canonical, $\theta=1$ in (20)), and deconvolution with our spatially varying hemodynamics (Corrected). We used scan-1 data from subject 103818 (hemodynamic field shown in Figure 2, left), with the seed ROI set to Left Ventral Area 6 (Glasser atlas) [43]. This ROI was chosen because it exhibited the largest average $\widehat{\theta}$ ($\approx 1.5$), thus providing strong contrast with the spatially constant model.
Figure 4 (left) shows the spatial maps of the $\log_{10}-$transformed FDR corrected p-values from the GC tests. To interpret this spatial pattern, we calculated the proportion of mesh vertices in each parcel of the Glasser atlas where the GC was significant at the $0.05$ level. Regions with $\geq 50\%$ of significant mesh vertices were deemed strongly effectively connected, and are shown in the right panel of Figure 4. We see that all of the strongly connected regions identified by BOLD and Canonical are also detected by our deconvolved signals (Corrected). However, the effective connectivity estimated under our corrected HRF identifies many additional regions, primarily located within the dorsal visual stream, occipital lobe and parietal lobe. This pattern is highly consistent with the established functional connectivity of Left Ventral Area 6 reported in the literature [44].
VI Conclusion and Discussion

### v2

@@ 5.2.3 Connectivity Analysis
Using the procedure in Section 4.5.2, we compared the seed-based effective connectivity for raw BOLD (BOLD), deconvolution with a spatially fixed canonical response (Canonical, $\theta=1$ in (20)), and deconvolution with our spatially varying hemodynamics (Corrected). We used scan-1 data from subject 103818 (hemodynamic field shown in Figure 3, left), with the seed ROI set to Left Ventral Area 6 (Glasser atlas) (Glasser et al., 2016). This ROI was chosen because it exhibited the largest average $\widehat{\theta}$ ($\approx 1.5$), thus providing strong contrast with the spatially constant model.
Figure 5 (left) shows the spatial maps of the $\log_{10}-$transformed FDR corrected p-values from the GC tests. To interpret this spatial pattern, we calculated the proportion of mesh vertices in each parcel of the Glasser atlas where the GC was significant at the $0.05$ level. Regions with $\geq 50\%$ of significant mesh vertices were deemed strongly effectively connected, and are shown in the right panel of Figure 5. We see that all of the strongly connected regions identified by BOLD and Canonical are also detected by our deconvolved signals (Corrected). However, the effective connectivity estimated under our corrected HRF identifies many additional regions, primarily located within the dorsal visual stream, occipital lobe and parietal lobe. This pattern is highly consistent with the established functional connectivity of Left Ventral Area 6 reported in the literature (Baker et al., 2018).
6 Conclusion and Discussion
@@
This work has a few limitation. While empirical results on both synthetic and in-vivo data indicate good calibration of the double bootstrap algorithm for uncertainty quantification, a limitation of this work is the computational burden of this procedure, as it requires $BR$ point estimates. In practice, we use $B=100$ and $R=20$ and ran the code in sequence. However, each of the outer bootstraps is independent and therefore can be parallelized, provided sufficient computational resources. In our experiments, we use a pairwise measure of Granger causality to estimate effective connectivity, which is somewhat unsatisfactory since it also detects indirect interactions. This choice was made for computational feasibility, as full multivariate Granger causality measures tend to scale poorly with the dimension of the times series (order $V^{2}$) (Shojaie and Fox, 2022), making application to the full surface-based mesh time series of dimension $V=32,492$ impractical.